# Scalable magnetic resonance fingerprinting: Incremental inference of high dimensional elliptical mixtures from large data volumes

### v1

@@ 7 Application to magnetic resonance fingerprinting (MRF) reconstruction
MRF is able to provide multiple quantitative tissue parameters images from shorter acquisition times, thanks to the simultaneous application of transient states excitation and highly undersampled $k$-space read-outs. These two aspects have a combined impact on acquisition times and image reconstruction accuracy. More undersampling allows more parameter estimations in reasonable acquisition times but is also responsible for larger undersampling errors, noise and artifacts, reducing map reconstruction accuracy.
In some earlier work (Oudoumanessah et al., 2024), the Gaussian version of our procedure, referred to as HD-GMM, was evaluated, for the reconstruction, in an ideal setting, of fully sampled acquisitions, targetting $L=3$ parameters. In this scenario, it was reported that HD-GMM, coupled with the online EM algorithm, achieved results comparable to full dictionary matching while significantly reducing reconstruction times. However, as we report in this section, HD-GMM performance degrades when dealing with largely undersampled acquisitions and becomes inadequate for estimating six parameters. In contrast, performance can be maintained by considering an elliptical version of our procedure (HD-MED) less sensitive to outliers.
@@
We compare two instances of HD-MED, namely HD-GMM and HD-STM, where the mixture components are respectively set to Gaussian and Student distributions. The computations for the corresponding OEM are detailed in Section 3 of the Appendix. All the experiments are performed with a Python code using the JAX library (Bradbury et al., 2018) with Nvidia V100-32gb GPU except for the dictionary generation part that is done with a mix of Matlab and Python. The Python code is available at https://github.com/geoffroyO/HD-MED.
@@ 7.1 Undersampled MRF acquisitions
In vivo acquisitions were conducted on 6 healthy volunteers (28$\pm$5.5 years old, 3 males and 3 females) using a 32-channel head receiver array on a Philips 3T Achieva dStream MRI at the IRMaGe facility (MAP-IRMaGe protocol, NCT05036629). This study was approved by the local medical ethics committee and informed consent was obtained from each volunteer prior to image acquisition. The imaging pulse sequence was based on an IR-bSSFP acquisition. 260 repetitions were acquired following the parameters proposed in Coudert et al. (2024). The acquisitions were performed using quadratic variable density spiral sampling (12 interleaves out of 13), matrix size=192x192x(4-5), voxel size=1.04x1.04x3.00 mm3 for a total scan duration of 2 minutes per slice. While the acquisition time may appear high for an MRF context compared to Gu et al. (2018), where full T1 and T2 maps are generated in 2 minutes, the longer sequences in Coudert et al. (2024) account for vascular parameters without requiring contrast agents.
Spiral sampling in MRI involves acquiring data in a spiral trajectory through k-space, covering the center first and gradually moving outward, which allows for faster data acquisition and more efficient use of the scanner’s gradients. However, this method leads to undersampling noise, manifesting as artifacts, because it captures less information about the image, reducing the ability to accurately reconstruct fine details (Körzdörfer et al., 2019). It results in a number $\tilde{N}_{s}\approx 140,000-180,000$ of in vivo MR signals per subject $s$, each of dimension $M=260$, and a total number $\tilde{N}$ of approximately $\tilde{N}=960,000$ signals to be matched for the reconstruction of $L=6$ parameter maps for 6 subjects.
@@ 7.2 MRF dictionary
As introduced, the MRF sequence used here is a bSSFP-derived sequence. This type of sequence has been preferred because of its sensitivity to local frequency distributions related to microvascular structures in the imaging voxel. To account for this, the MRF dictionary is generated using the approach described in Coudert et al. (2024). First a base dictionary ${\cal D}_{f_{0}}$ is simulated using Bloch-equations for different combinations of ($T_{1}$, $T_{2}$, $B_{1}$, $\delta\!f$) browsed through a 4-dimensional regular grid. Simulations were made at a magnetic field strength of $3.0$T, on a regular parameter grid made of $20$ $T_{1}$ values (from $200$ to $3500ms$), $20$ $T_{2}$ values (from $10$ to $600ms$), $10$ $B_{1}$ values (from $0.7$ to $1.2$) and $100$ frequency offset $\delta\!f$ values (from $-50$ to $49$ Hz with an increment of 1 Hz), keeping only signals for which T1$>$T2, resulting in an initial $390,000$ entries dictionary. These Bloch simulations are obtained using an in-house mix of Python and Matlab code based on initial implementation by B. Hargreaves (Hargreaves, ). The relaxometric parameters vary within the dictionary to allow their estimation. $B_{1}$ is varied to ensure the realism of the dictionary, given the sequence’s sensitivity to this parameter, which could otherwise bias the estimation of the other parameters. Finally, varying $\delta\!f$ values are necessary to compute microvascular contributions, as detailed below.
To build signals that capture additional microvascular network’s blood volume (CBV) and mean vessel radius (R) information, we follow a construction proposed by Coudert et al. (2024). Microvascular network segmentations are used to pre-compute 2500 different intra-voxel frequency distributions centred at $\delta\!f$ values. A CBV and R value characterizes each distribution. Using these pre-computed distributions, new signals are obtained by summing signals from the base dictionary weighted according to each frequency distribution. It results an expanded 6-dimensional dictionary ${\cal D}_{f}$ of almost $N=400,000,000$ signals of dimension $M=260$, encoding for ($T_{1}$, $T_{2}$, $B_{1}$, $\delta\!f$, $CBV$, $R$). Figure 2 gives an illustration on how this operation changes a given signal in the base dictionary into 2 new signals depending on values of $CBV$ and $R$.
Figure 2: Illustration of the base dictionary expansion. The blue signal corresponding to a non-blood voxel with associated $T_{1}$, $T_{2}$, $B_{1}$ and $\delta\!f$ values. By convoluting this signal along the $\delta\!f$ dimension with two different frequency distributions corresponding to different values of $CBV$ and $R$, the resulting orange and green signals are then obtained for voxels with the same $T_{1}$, $T_{2}$, $B_{1}$ and $\delta\!f$ values as the blue signal but with different $CBV$ and $R$ values.
@@ 7.3 Model selection and initialization
The only two hyperparameters that need to be set are the number of mixture components $K$ and the desired reduced dimension vector ${\mathb{d}}$. For mixture models, $K$ can be selected using the Bayesian Information Criterion (BIC) (Schwarz, 1978), which requires running multiple models with varying $K$ values and finding the elbow or the minimum of the BIC curve. The desired dimension vector ${\mathb{d}}$ is automatically set during an initialization step, which makes $K$ the only parameter that the user needs to set prior to running the OEM algorithm.
The initialization step includes determining the reduced dimension vector ${\mathb{d}}$. We randomly choose a subset of the dictionary that fits into memory and run a batch full covariance EM algorithm to determine a first estimation of the ${\boldsymbol{\Sigma}}_{k}$’s. These estimations are decomposed into eigenvalues and eigenvectors to determine ${\mathb{D}}_{k}$’s and ${\mathb{A}}_{k}$’s. For each $k$, $d_{k}$ is then determined by applying a scree plot to the eigenvalues using the kneedle algorithm (Satopaa et al., 2011), ${\mathb{a}}_{k}$ is then initialized to the first $d_{k}$ eigenvalues, and the last eigenvalues are averaged to initialize $b_{k}$. This kind of initialization of ${\mathb{D}}_{k}$ and ${\mathb{A}}_{k}$ is called spectral initialization and is similar to the one proposed by Hong et al. (2021), proving to be relatively stable compared to other types of initializations.
In Figure 3 (left), we show the different BIC curves for HD-GMM and HD-STM models trained on the dictionary of signals with $K$ varying from $5$ to $80$. The elbows of the curves are found at $K=30$ for HD-GMM, and at $K=25$ for HD-STM. Figures 3 (middle and right) show the different vectors ${\mathb{d}}$ obtained for varying $K$. The larger the points, the higher the number of components having the reduced dimension indicated on the $y$-axis. Most of these dimensions are between $5$ and $40$, meaning that the original $M=260$ signal dimension can be reduced by a factor of approximately $7$ to $10$ leading to a reduction from $13$ To to approximately $1.5$ To of signals stored in memory. To check, how much is lost in this reduction, we compute the root mean square error (RMSE) between the dictionary $\mathbf{y}_{i}$ signals and their reconstructions $\check{\mathbf{y}}_{i}$ from their reduced representations (11),
$RMSE=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(\mathbf{y}_{i}-\check{\mathbf{y}}_{i})^{2}}\;.$
@@
Figure 5 shows as expected that the reconstruction RMSE decreases when $K$ increases and is smaller for HD-STM in particular for small values of $K$. For the selected $K$ values it represents about 2.5% of the signal.
Figure 4: Dictionary reduction loss as measured by RMSE values over the dictionary signals, for HD-GMM and HD-STM, with respect to the number of components $K$.
Figure 5: Parameter maps comparison. Average SSIM in [-1,1], the higher the better, and standard deviation over subjects, between HD-GMM (resp. HD-STM) and full matching maps.
@@ 7.4 Tissue and sensitivity parameters reconstruction
For the six subjects, we compare the parameter maps obtained with our Algorithm 1 using HD-GMM and HD-STM to that obtained with traditional matching, referred to as full matching (i.e., matching with uncompressed signals). While full matching cannot be considered the ground truth, it remains the reference method due to its robustness to the slightly undersampled MRF acquisitions used in this study, as demonstrated in (Coudert et al., 2024), despite its high computational and memory demands.
Table 1 reports for each of the 6 parameters, the distances to full matching parameter estimations, showing the average Mean Absolute Errors ($MAE$) across voxels for all $n_{subject}=6$ subjects and all slices. For another comparison that takes into account the image structure, Figure 5 shows the reconstruction quality as measured by the structural similarity index measure (SSIM) (Wang et al., 2004), for both HD-STM and HD-GMM when compare to full matching, for each parameter and averaged over subjects. The SSIM is a decimal value between -1 and 1, where 1 indicates perfect similarity, 0 indicates no similarity, and -1 indicates perfect anti-correlation.
@@
Table 1: Parameter maps reconstruction. Average MAE and standard deviation, over voxels and subjects, for HD-GMM ($K=30$) and HD-STM ($K=25$) with respect to full matching. Best values in bold.
[…]
@@
HD-STM consistently outperforms HD-GMM across all parameters, as evidenced by Figures 5 and Table 1. HD-STM achieves superior SSIM and MAE values, indicating better structural correspondence. This advantage is particularly pronounced for parameters such as $T_{1}$ and $T_{2}$, where HD-STM demonstrates robust reconstruction capabilities, whereas HD-GMM fails to accomodate too much undersampling noise. However, for parameters like CBV and $R$, HD-STM’s performance declines, suggesting that these parameters are inherently more challenging to model accurately. In these cases, HD-GMM yields the lowest SSIM values and the highest MAE values, further underscoring its limitations. Overall, while HD-STM proves to be the more reliable method, the significant variability across subjects—particularly for $\delta\!f$ and CBV—and the observed SSIM drop for $R$ highlight the need for further refinements.
@@
For another assessment of the maps quality, Table 2 presents the mean values and standard deviations of $T_{1}$, $T_{2}$, $CBV$, and $R$, computed over voxels in white and gray matter ROIs delineated on $T_{1}$ maps obtained from full matching. Compared to the ranges for healthy subjects reported by Wansapura et al. (1999); Bjørnerud and Emblem (2010); Delphin et al. (2023), HD-STM produces values that are both more consistent with expected ranges and closer to those obtained using full matching than to those derived from HD-GMM. In particular, with HD-GMM, $T_{2}$ values significantly depart from the literature reference, as also visible in Figure 6.
@@
Indeed, the conclusions drawn from Figure 5 and Tables 12 are further illustrated in Figure 6. This figure highlights HD-GMM’s failure to reconstruct parameters accurately, even when some estimates fall within healthy ranges reported in the literature from Table 2. In contrast, HD-STM consistently delivers parameter estimates for $T_{1}$, $T_{2}$, $\delta\!f$, and $B_{1}$ that are nearly artifact-free and closely aligned with expected values. Additionally, HD-STM effectively captures primary structures and generates homogeneous maps for vascular parameters such as $CBV$ and $R$ even though minor residual shim artifacts persist, these are likely due to the high degree of undersampling in the acquired images.
@@
Figure 6: $T_{1}$, $T_{2}$, $\delta\!f$, $B_{1}$ sensitivity, $CBV$, $R$ maps (columns) with different methods (lines) - full matching, HD-STM ($K=25$) and HD-GMM ($K=30$). Results for all other slices of this subject are shown in Appendix Section 5.
@@
Table 2: Mean $T_{1},T_{2},\text{CBV},R$ values and standard deviations, in white (WM) and grey (GM) matters ROIs. Values significantly departing from literature reference (last column) are in red.
Parameter ROI Full matching HD-GMM HD-STM Literature (Wang et al., 2019b)
$T_{1}$(ms) WM $891\pm 6$ $650\pm 23$ $869\pm 19$ $690$-$1100$
GM $1566\pm 12$ $1650\pm 57$ $1631\pm 56$ 1286-1393
$T_{2}$(ms) WM $52\pm 0.0$ $400\pm 15$ $48\pm 2$ 56-80
GM $95\pm 4$ $370\pm 43$ $96\pm 12$ 78-117
[…]
GM $9.2\pm 1.1$ $4.2\pm 0.8$ $7.6\pm 1$ $3.0$ $8.0$
$R$ ($\mu s$) WM $5.7\pm 0.1$ $5.6\pm 0.1$ $5.7\pm 0.1$ $6.8\pm 0.3$
GM $6.2\pm 0.1$ $5.4\pm 0.4$ $6.1\pm 0.1$ $7.3\pm 0.3$

### v2

@@ 7 Application to magnetic resonance fingerprinting (MRF) reconstruction
MRF provides multiple tissue parameter maps from shorter acquisition times, thanks to the simultaneous application of transient states excitation and highly undersampled $k$-space read-outs. These two aspects have a combined impact on acquisition times and image reconstruction accuracy. More undersampling allows more parameter estimations but is also responsible for larger undersampling errors, noise and artifacts, reducing map reconstruction accuracy.
In some earlier work (Oudoumanessah2025), the Gaussian version of our procedure, referred to as HD-GMM, was evaluated, for the reconstruction, in an ideal setting, of fully sampled acquisitions, targeting $L=3$ parameters. In this scenario, it was reported that HD-GMM, coupled with the online EM algorithm, achieved results comparable to full dictionary matching while significantly reducing reconstruction times. However, as we report in this section, HD-GMM performance degrades when dealing with largely undersampled acquisitions and becomes inadequate for estimating six parameters. Performance can be maintained by considering an elliptical version of our procedure (HD-MED) less sensitive to outliers.
@@
We compare two instances of HD-MED, namely HD-GMM and HD-STM, where the mixture components are respectively set to Gaussian and Student distributions. The computations for the corresponding OEM are detailed in Section 5 of the Appendix. Additional computations and results for mixtures of Laplace distributions (HD-LM) are given in Appendix Section 6. All experiments are performed with a Python code using the JAX library (jax2018github) with Nvidia H100-80gb GPU for training, and V100-32gb GPU for inference except for the dictionary generation part that is done with a mix of Matlab and Python.
@@ 7.1 Undersampled MRF acquisitions
[…]
Spiral sampling in MRI involves acquiring data in a spiral trajectory through k-space, covering the center first and gradually moving outward, which allows for faster data acquisition and more efficient use of the scanner’s gradients. However, this method leads to undersampling noise, manifesting as artifacts, because it captures less information about the image, reducing the ability to accurately reconstruct fine details (korzdorfer2019). It results in a number $\tilde{N}_{s}\approx 140,000-180,000$ of in vivo MR signals per subject $s$, each of dimension $M=260$, and a total number $\tilde{N}$ of approximately $\tilde{N}=960,000$ signals to be matched for the reconstruction of $L=6$ parameter maps for 6 subjects.
@@ 7.2 MRF dictionary
The MRF sequence used here is a bSSFP-derived sequence. This type of sequence has been preferred because of its sensitivity to local frequency distributions related to microvascular structures in the imaging voxel. The MRF dictionary is generated using the approach described in coudert2024. First, a base dictionary ${\cal D}_{f_{0}}$ is simulated using Bloch equations for different combinations of ($T_{1}$, $T_{2}$, $B_{1}$, $\delta\!f$) sampled over a 4-dimensional regular grid. Simulations were made at a magnetic field strength of 3.03.0T, on a regular parameter grid made of 2020 $T_{1}$ values (from $200$ to $3500ms$), 2020 $T_{2}$ values (from $10$ to $600ms$), $10$ $B_{1}$ values (from $0.7$ to $1.2$) and $100$ frequency offset $\delta\!f$ values (from 50-50 to $49$ Hz with an increment of 1 Hz), keeping only signals for which T1>>T2, resulting in an initial $390,000$ entries dictionary. These Bloch simulations are obtained using an in-house mix of Python and Matlab code based on initial implementation by B. Hargreaves (Bloch_Simulator). The relaxometric parameters vary within the dictionary to allow their estimation. $B_{1}$ is varied to ensure the realism of the dictionary, given the sequence’s sensitivity to this parameter, which could otherwise bias the estimation of the other parameters. Finally, varying $\delta\!f$ values are necessary to compute microvascular contributions, as detailed below.
To build signals that capture additional microvascular network’s blood volume (CBV) and mean vessel radius (R) information, we follow the construction of coudert2024. Microvascular network segmentations are used to pre-compute 2500 different intra-voxel frequency distributions centred at $\delta\!f$ values. A CBV and R value characterize each distribution. Using these pre-computed distributions, new signals are obtained by summing signals from the base dictionary weighted according to each frequency distribution. The result is an expanded 6-dimensional dictionary ${\cal D}_{f}$ of almost $N=400,000,000$ signals of dimension $M=260$, encoding for ($T_{1}$, $T_{2}$, $B_{1}$, $\delta\!f$, $CBV$, $R$). Figure 2 illustrates how this operation changes a signal in the base dictionary into 2 new signals depending on $CBV$ and $R$ values.
Figure 2: Illustration of the base dictionary expansion. The blue signal corresponding to a non-blood voxel with associated $T_{1}$, $T_{2}$, $B_{1}$ and $\delta\!f$ values. By convoluting this signal along the $\delta\!f$ dimension with two different frequency distributions corresponding to different values of $CBV$ and $R$, the resulting orange and green signals are then obtained for voxels with the same $T_{1}$, $T_{2}$, $B_{1}$ and $\delta\!f$ values as the blue signal but with different CBVCBV and $R$ values.
@@ 7.3 Model selection and initialization
In practice, the online EM algorithm is run in batches (Nguyen:2020ac) of size $B=2048$ signals at each iteration $r$ where $r\in[1:R]$ for typically $R=\lceil N/B\rceil$. When $N$ is huge, the $N$ signals are read only once in a single epoch, while for more moderate $N$, several epochs, e.g. 10, can be run with shuffling, making use of the fact that samples are available and can be revisited. Regarding the learning rates, $(\gamma_{r})_{r=1}^{R}$, they are set to $\gamma_{r}=(1-\epsilon)r^{-0.5-\epsilon^{\prime}}$ with $\epsilon=10^{-10},\epsilon^{\prime}=10^{-1}$ and Polyak-Ruppert averaging, following theoretical recommendations. In particular, for this choice, the MSE converges at the optimal rate and the covariance in the CLT is minimal, see Lauand2024 for additional results and a recent literature survey.
The only two other hyperparameters that need to be set are the number $K$ of components and the desired reduced dimension vector $\mathb{d}$}. Once initialized, these quantities are considered as fixed and not updated during the online procedure. For mixture models, $K$ can be selected using the Bayesian Information Criterion (BIC) (schwarz1978), which requires running multiple models with varying $K$ values and finding the elbow or the minimum of the BIC curve. The desired dimension vector $\mathb{d}$} is automatically set during an initialization step, which makes $K$ the only parameter that the user needs to set prior to running the OEM algorithm.
The initialization step includes determining the reduced dimension vector $\mathb{d}$}. We randomly choose a subset of the dictionary that fits into memory and run a batch full covariance EM algorithm to determine a first estimation of the ${\boldsymbol{\Sigma}}_{k}$’s. In practice, a subset of 10,000 signals was used. The covariance estimations are then decomposed into eigenvalues and eigenvectors to determine ${\mathb{D}}_{k}$’s and ${\mathb{A}}_{k}$’s. For each $k$, $d_{k}$ is then determined by applying a scree plot to the eigenvalues using the kneedle algorithm (satopaa2011), ${\mathb{a}}_{k}$ is then initialized to the first $d_{k}$ eigenvalues, and the last eigenvalues are averaged to initialize $b_{k}$. This kind of initialization of ${\mathb{D}}_{k}$ and ${\mathb{A}}_{k}$ is called spectral initialization and is similar to the one proposed by hong2021, proving to be relatively stable compared to other types of initializations.
In Figure 3 (left), we show the different BIC curves for HD-GMM and HD-STM models trained on the dictionary of signals with $K$ varying from $5$ to $80$. The elbows of the curves are found at $K=30$ for HD-GMM, and at $K=25$ for HD-STM. Figures 3 (middle and right) show the different vectors $\mathb{d}$} obtained for varying $K$. The larger the points, the higher the number of components having the reduced dimension indicated on the $y$-axis. Most of these dimensions are between $5$ and $40$, meaning that the original $M=260$ signal dimension can be reduced by a factor of approximately $7$ to $10$ leading to a reduction from $13$ TB to approximately $1.5$ TB of signals stored in memory. To check, how much is lost in this reduction, we compute the root mean square error (RMSE) between the dictionary $\mathbf{y}_{i}$ signals and their reconstructions $\check{\mathbf{y}}_{i}$ from their reduced representations (11), $RMSE=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(\mathbf{y}_{i}-\check{\mathbf{y}}_{i})^{2}}\;.$
@@
Figure 5 shows that the RMSE decreases when $K$ increases and is smaller for HD-STM in particular for small $K$. For the selected $K$ values it represents about 2.5% of the signal.
Figure 4: Dictionary reduction loss measured by RMSE over the dictionary signals, for HD-GMM and HD-STM, with respect to the number of components $K$.
Figure 5: Parameter maps comparison. Average SSIM in [-1,1], the higher the better, and standard deviation over subjects, between HD-GMM, (resp. HD-STM, DRONE) and full matching maps.
Similar plots are shown in Figure 1 of Appendix Section 6 for the Laplace model HD-LM, leading to select $K=30$ and similar cluster dimension values.
@@ 7.4 Tissue and sensitivity parameters reconstruction
For the six subjects, we compare the parameter maps obtained with our Algorithm 1 using HD-GMM, HD-STM and with a MRF state-of-the-art neural network approach DRONE (cohen2018) tailored for $T_{1}$ and $T_{2}$, which we adapted for additional parameters, to that obtained with traditional matching, referred to as full matching (i.e., matching with uncompressed signals). While full matching cannot be considered as the ground truth, it remains the reference method due to its robustness to the highly undersampled MRF acquisitions used in this study, as demonstrated in (coudert2024), despite its high computational and memory demands. Conventionally, the distance to compare two signals for matching is the Euclidean distance.
Table 1 reports for each of the 6 parameters, the distances to full matching parameter estimations, showing the average Mean Absolute Errors ($MAE$) across voxels for all $n_{subject}=6$ subjects and all slices. For another comparison that takes into account the image structure, Figure 5 shows the reconstruction quality as measured by the structural similarity index measure (SSIM) (Wang2004) when compared to full matching, for each parameter and averaged across subjects. The SSIM is a decimal value between -1 and 1, where 1 indicates perfect similarity, 0 indicates no similarity, and -1 indicates perfect anti-correlation.
@@
Table 1: Parameter maps reconstruction. Average MAE and standard deviation, over voxels and subjects, for HD-GMM ($K=30$) HD-STM ($K=25$) and DRONE with respect to full matching. Best values in bold.
Parameter $T_{1}$ (ms) $T_{2}$ (ms) $\delta f$ (Hz) $B_{1}$ sensitivity ($10^{-3}$) CBV ($\%$) $R$ ($\mu m$)
@@
HD-STM consistently outperforms HD-GMM and DRONE for all parameters, as evidenced by Figures 5 and Table 1. HD-STM achieves superior SSIM and MAE values, indicating better structural correspondence. This advantage is particularly pronounced for parameters such as $T_{1}$ and $T_{2}$, where HD-STM demonstrates robust reconstruction capabilities, whereas HD-GMM, and to a lesser extent DRONE, fail to accommodate too much undersampling noise. However, for parameters like CBV and $R$, HD-STM’s performance decreases, suggesting that these parameters are inherently more challenging to model accurately. In these cases, DRONE and HD-GMM yield the lowest SSIM values and the highest MAE values, further underscoring their limitations. Overall, while HD-STM proves to be the more reliable method, the significant variability across subjects—particularly for $\delta\!f$ and CBV—and the observed SSIM drop for $R$ highlight the need for further refinements.
@@
For another quality assessment, Table 2 presents the mean and standard deviations of $T_{1}$, $T_{2}$, $CBV$, and $R$, computed over voxels in white and grey matter ROIs delineated on $T_{1}$ maps obtained from full matching. Compared to the ranges for healthy subjects reported by wansapura1999; bjornerud2010; delphin2023, HD-STM produces values that are both more consistent with expected ranges and closer to those obtained using full matching than to those derived from HD-GMM or DRONE. In particular, with HD-GMM (resp. DRONE), $T_{2}$ (resp. $R$) values significantly depart from the literature reference, as also visible in Figure 6.
@@
The conclusions drawn from Figure 5 and Tables 12 are further illustrated in Figure 6. This figure highlights HD-GMM and DRONE’s failure to reconstruct parameters accurately, even when some estimates fall within healthy ranges reported in the literature from Table 2. In contrast, HD-STM consistently delivers parameter estimates for $T_{1}$, $T_{2}$, $\delta\!f$, and $B_{1}$ that are nearly artifact-free and closely aligned with expected values. Additionally, HD-STM effectively captures primary structures and generates homogeneous maps for vascular parameters such as $CBV$ and $R$ even though minor residual shim artifacts remain, likely resulting from significant undersampling in the acquired images. Nevertheless, the presence of these artifacts indicates that HD-STM preserves the physical characteristics of the image, rather than suppressing them. Reconstruction results for HD-LM are shown in Appendix Section 6.1. HD-LM performs much better than HD-GMM and DRONE but worse than HD-STM.
@@
Figure 6: $T_{1}$, $T_{2}$, $\delta\!f$, $B_{1}$ sensitivity, CBVCBV, $R$ maps (columns) with different methods (lines) - full matching, HD-STM ($K=25$), HD-GMM ($K=30$), and DRONE. All other slices of this subject are shown in Appendix Section 8, Figure 3.
@@
Table 2: Mean $T_{1},T_{2},\text{CBV},R$ and standard deviations, in white (WM) and grey (GM) matter ROIs. Values significantly departing from literature (last column) are in red. 1: cohen2018.2: WangT2etoile
Parameter ROI Full matching HD-GMM HD-STM DRONE111 Literature222
$T_{1}$(ms) WM $891\pm 6$ $650\pm 23$ $869\pm 19$ $696\pm 14$ 690690-$1100$
GM $1566\pm 12$ $1650\pm 57$ $1631\pm 56$ $1660\pm 53$ 1286-1393
$T_{2}$(ms) WM $52\pm 0.0$ $400\pm 15$ $48\pm 2$ $125\pm 5$ 56-80
GM $95\pm 4$ $370\pm 43$ $96\pm 12$ $135\pm 24$ 78-117
CBV($\%$) WM $5.4\pm 0.6$ $2.0\pm 0.1$ $4.8\pm 0.8$ $11.5\pm 0.3$ $1.7$ - $3.6$
GM $9.2\pm 1.1$ $4.2\pm 0.8$ $7.6\pm 1$ $14.6\pm 0.6$ $3.0$ - $8.0$
$R$ ($\mu s$) WM $5.7\pm 0.1$ $5.6\pm 0.1$ $5.7\pm 0.1$ $0.5\pm 0.01$ $6.8\pm 0.3$
GM $6.2\pm 0.1$ $5.4\pm 0.4$ $6.1\pm 0.1$ $0.6\pm 0.1$ $7.3\pm 0.3$